What is the simplified form of (-9)^1/2
Answer by amoresroy(361)

What is the simplified form of (-9)^1/2
= (9)^1/2 * (-1)^1/2
= 3 * (-1)^1/2
= 3i
note that (-1)^1/2 = i (square root of negative 1)